# Correcting Nonresponse Bias Using Panel Data

### v3

@@ Abstract
When subjects who respond to requests for data, such as in surveys or post-treatment follow-up, are not representative of the population as a whole, inferences drawn from the data can be misleading. We show that if subjects’ accumulated requests and responses over time are recorded and organized as panel data, requests can be used as instruments to correct for nonresponse bias even if total requests are not randomized between subjects. We demonstrate our method by estimating an 18-percentage-point gender gap in entrepreneurial career intentions using a survey of undergraduates at the University of Wisconsin-Madison with a 15% response rate and a 20-percentage-point intention gap among respondents.
@@ 1 Introduction
In Section 3, we demonstrate the use of nonrandom request instruments by estimating the gender gap in entrepreneurial career intentions among undergraduates using survey data from the University of Wisconsin-Madison. Estimating a parametric van1981demand model for binary outcomes, we find an 18-percentage-point nonresponse-corrected male-female gap in entrepreneurial intentions. We find statistically insignificant evidence of positive selection for men and negative selection for women driving the 20 percentage point uncorrected gap among respondents. This model is identified with two requests, so we also implement and pass a “predictable trends” overidentification test of the null that later requests have no direct effect on the outcome, supporting the validity of our parametric and nonparametric assumptions in this application. In Appendix B, we estimate labor market outcomes using a synthetic version of the Norway in Corona Times survey. We construct our synthetic dataset to match respondent mean estimates reported by dutz2021selection, which are sufficient to estimate population means with our method. The 95% confidence intervals of our estimated population means cover ground truth averages from administrative data reported by these authors for five out of six survey variables, with respondent averages differing significantly from the ground truth for all variables.
@@ 2.2 Modeling Nonresponse
We assume that observed variables are realizations of potential outcomes, following rubin1974estimating. First, let $S_{it}=S_{it}(R_{it})\prod_{j=0}^{t-1}(1-S_{ij}(R_{ij}))$ where $S_{it}(r)\in\{0,1\}$ indicates subject $i$’s willingness to respond to $r$ requests at time tt. [^7] We use this definition in lieu of $S_{it}=S_{it}(R_{it})$ because we assume that subjects respond at most once. Similarly, let $Y_{it}=Y_{it}(S_{it},R_{it})=Y_{it}(1,R_{it})S_{it}+Y_{it}(0,R_{it})(1-S_{it})$ with $Y_{it}(s,r)$ denoting the response subject $i$ would give at time $t$ if they made response choice $s$ given $r$ accumulated requests, with $Y_{it}(0,r)$ missing for all $i$, $t$, and $r$. $S_{it}(r)$ and $Y_{it}(s,r)$ are defined for all values in the supports of their domains, $r\in\{0,...,\max(R_{it})\}$ and $s\in\{0,1\}$, respectively, while they are observed only for their realized values.
@@
It follows that we can learn about $m(u)$ if we can estimate $\mathbb{E}[Y_{i}^{*}|S_{it}(r)-S_{it}(r^{\prime})=1]$, which we refer to as the local average response for $r-r^{\prime}$ compliers. If $m(u)$ is parameterized, as in a heckman1979sample model, point estimation is possible if there are weakly more identified values of $P(r)$ than parameters. Otherwise, $m(u)$ can be bounded using methods such as those of horowitz2000nonparametric or lee2009training. Local average responses are identified as
@@ 2.3 Local Average Response Identification Assumptions
[…]
@@
The importance of avoiding experimenter demand effects or priming in surveys is well established (stantcheva2023run). Our assumption expresses this as a special case of similar assumptions made in the context of sample selection by heckman1979sample and in the context of treatment effect identification by angrist1996identification. In contrast with intertemporal independence, exclusion is more likely to be violated due to continuous time variation in potential responses and requests if time periods are defined coarsely, as subjects receiving different numbers of requests in the same discrete time period may receive them at different points in continuous time. [^9] Following imbens1994identification, we’ve assumed implicitly that $Y_{it}(s,r)$ and $S_{it}(r)$ are fixed for all $s,r$, for all $i,t$ (SUTVA). This assumption is consequentially violated if potential responses or potential response choices systematically vary in continuous time within discrete time periods.
@@
This assumption is weaker than $Y_{it}(1,r)=Y_{i}^{*}$, which is common in the related literature, because it allows for idiosyncratic shocks that cancel out within complier groups, such as if students’ potential responses regarding career intentions change over time in response to new information, such as grades on assignments. [^10] We hope that our consideration of measurement error will forestall inaccurate presumptions that weaker assumptions, such as $\mathbb{E}[Y_{it}(1,r)]=\mathbb{E}[Y_{i}^{*}]$, are sufficient. Intertemporal independence and exclusion rule out systematic time-varying measurement error and measurement error caused by requests, respectively, but they do not rule out systematic inaccuracies within or across complier groups. For example, if less conscientious subjects give systematically inaccurate responses to late requests, nonresponse corrections using request instruments will tend to replicate similar inaccuracy in their predictions for non-respondents.
@@ 3 Application: Gender Entrepreneurship Gap
We use our method to estimate the gender gap in entrepreneurial aspirations among undergraduate students. There is a large gender gap in entrepreneurship among working-age adults (aldrich2005entrepreneurship), perhaps driven at least in part by a gender gap in early-stage funding for new ventures (canning2012women; greene2003women). Our paper contributes to this literature by investigating pre-market gender gaps in entrepreneurial career interest, which may be more likely to develop due to intrinsic interest gaps or pre-market discrimination rather than labor market discrimination. This application showcases the importance of researcher judgments in defining the time index, adjusting for nonrandom variation in requests, and testing modeling assumptions in cases where there are more requests than are needed for identification. In Appendix B, we demonstrate our method’s accuracy by showing that nonresponse-corrected estimates are closer to ground truth values from administrative data than respondent averages for the Norway in Corona Times Survey, using survey and ground truth averages reported by dutz2021selection.
@@ 3.1 Data
We use a survey of the entrepreneurial intentions of the undergraduate population at University of Wisconsin–Madison that was implemented every fall from 2015 to 2022, as well as in spring 2020 and 2021. Students could choose between “Yes,” “No,” or “I Don’t Know,” when responding to a question regarding whether they intend to pursue a career in entrepreneurship. We code an “Intention” variable equal to 1 if the student answered “Yes” and 0 if they answered “No” or “I Don’t Know.” We observe term-specific survey request stratum indicators for groups of students who are sent requests at the same time through survey software. We observe timestamps, in seconds, for each request made to each stratum and for each student’s response if they respond. Students who opted out of the survey in a prior year have no stratum indicators and no request or response timestamps in the raw data, whereas students who respond to early requests have stratum indicators and request timestamps for all intended requests in raw data, though they did not receive additional actual requests after responding.
@@ 3.2 Constructing a Request-Response Panel
Our results from Section 2 require panel data wherein subjects are observed multiple times, so we construct such a panel. In our setting, students who opted out of the survey in a prior term and students who respond early in a given term do not receive subsequent requests. It follows that actual requests almost surely violate the independence assumption in Section 2, as students receiving each number of requests differ systematically. We address this by randomly assigning opt-outs to strata ex post and assuming $S_{i}(r)=0$ for all $r$ for these students, then imputing intended request timestamps for each student using the request timestamps of nonrespondents in their stratum.
For each student-term in a stratum, we produce observations for $t$ in $0,1,...,R_{i}$ such that time periods are defined as intervals in which a subject has received a given number of intended requests, with $R_{it}=t$ for all $i$ and $t$. Independence and monotonicity require that the times at which requests were made had representative potential responses and that subjects had sufficient time to respond between requests, respectively. We report request timestamps in Appendix A, revealing that the shortest interval between requests is three days, which we assume was adequate time for students to respond. We set $S_{it}=1$ if subject $i$’s response timestamp is between those defining the beginning and end of period $t$, with $Y_{it}$ given by their response to the survey if $S_{it}=1$. We forward-impute retained response choices $\hat{S}_{it}$ and retained responses $\hat{Y}_{it}$ as described in Section 2.
@@ 3.3 Nonresponse-Corrected Entrepreneurial Intentions by Gender
We begin by estimating the raw gender gap in entrepreneurial intentions without adjusting for the covariates listed in Table 1. Because requests vary in number from two to seven between terms, we drop all observations with $R_{it}>2$ to support our independence assumption for the current specification. We estimate a parametric van1981demand selection model separately by gender via maximum likelihood, in which
$\hat{S}_{it}=\begin{cases}&1\mbox{ if }Z_{it}\alpha+u_{i}\geq 0,\\ &0\mbox{ otherwise,}\end{cases}$ (4)
$\hat{Y}_{it}=\begin{cases}1\mbox{ if }X_{i}\beta+\epsilon_{i}\geq 0\quad\&\quad&\hat{S}_{it}=1,\\ 0\mbox{ if }X_{i}\beta+\epsilon_{i}<0\quad\&\quad&\hat{S}_{it}=1,\\ \cdot\mbox{ if }&\hat{S}_{it}=0,\end{cases}$ (5)
@@
where $X_{i}=1$, $Z_{it}=[X_{i},I(R_{it}=2)]$, and $\rho$ gives the correlation between outcome residuals $\epsilon_{i}$ and response preference $u_{i}=\Phi^{-1}(1-U_{i})$, where $U_{i}$ is $i$’s response aversion percentile in the notation of Section 2 and $\Phi^{-1}(\cdot)$ denotes the inverse of the normal CDF. [^12] Our 15% response rate is insufficient for informative bounds from conservative approaches such as that of horowitz2000nonparametric, and we doubt that gender satisfies the monotonicity assumption of lee2009training or behaghel2015please, rendering the “local gender gap” among survey respondents unidentified, to say nothing of its policy-relevance.
This model implies that expected intention as a function of response aversion is $m(u)=\Phi((X_{i}\beta+\rho\Phi^{-1}(1-u))/\sqrt{1-\rho^{2}})$, with the expectation conditional only on $X_{i}$ given by $\int_{0}^{1}m(u)du=\Phi(X_{i}\beta)$. We graphically present our estimates of $m(u)$ for men and women in Figure 1 for $u\in[0,1]$. Local average responses and sample averages for compliers to the first two requests are shown for reference. The levels of entrepreneurial intention for early and late respondents show a decreasing pattern for men and an increasing pattern for women, as in Table 1. Differences in selection bias between men and women, reflected by their $\rho$ estimates, explain the difference between the 20 percentage point uncorrected gap estimate and the 12 percentage point corrected gap estimate.
@@ 3.4 Oaxaca-Blinder-Kitagawa Decomposition
To implement the decomposition, we estimate the model in (4)-(6) jointly for men and women via maximum likelihood using gender-specific parameters $(\beta_{m},\alpha_{m},\rho_{m})$ for men and $(\beta_{w},\alpha_{w},\rho_{w})$ for women. To avoid overweighting terms in which more requests were sent, which have more observations, we weight each observation by the inverse of the total number of requests sent to that student in that term. We define $X_{i}$ as all covariates listed in Table 1 along with term fixed effects, and we define $Z_{it}$ as binary indicators for each value in $R_{it}=1,2...,7$ interacted with $X_{i}$.131313blandhol2022tsls recommend rich covariate interactions for instrumental variables when estimating local average treatment effects via two stage least squares. The estimator and target parameters are different here, but we find their argument convincing that if model assumptions hold conditional on $X_{i}$, then $X_{i}$ should be flexibly interacted with instruments to support identification. We order the sample of $N_{w}$ women and $N_{m}$ men such that subjects with $i=1,...,N_{w}$ are women and those with $i=N_{w}+1,...,N$ are men, with $N=N_{w}+N_{m}$.
@@
Notes: Uncorrected estimates obtained by estimating a van1981demand model using only the final request observation for each student-term with the constraint $\rho=0$. Corrected estimates obtained by estimating the same model using all student-term-request observations with unconstrained $\rho$, with observations weighted by the inverse of total requests by student-term. Group-specific population mean estimates of variables and parameters, as indicated, are in columns (1), (2), (4), and (5), with sample means of $X$ calculated across student-term observations. Columns (3) and (6) give uncorrected and corrected estimated contributions of each variable or coefficient to the estimated gap in entrepreneurial intentions, using equations (7), (8), (9), (10), and (11). $\mathbb{E}[\mathbb{E}[Y_{i}^{*}|S_{i}=1,X_{i}]]$ is estimated as the average of $\Phi(X_{i}\hat{\beta}^{uncorrected})$ and $\mathbb{E}[\mathbb{E}[Y_{i}^{*}|X_{i}]]$ is estimated as the average of $\Phi(X_{i}\hat{\beta}^{corrected})$ over student-term observations. Term fixed effects, included in $X_{i}$ to support the independence assumption, are treated as part of the unexplained gap. Standard errors clustered at the individual level are in parentheses.
We estimate the contribution of $x_{i}\subset X_{i}$ to the entrepreneurial intention gap as
$\hat{\Delta}_{x}=\sum_{i=1}^{N_{w}}\frac{\Phi(X_{i}\hat{\beta}_{w}+(\bar{x}_{m}-\bar{x}_{w})\hat{\beta}_{w,x})-\Phi(X_{i}\hat{\beta}_{w})}{N_{w}},$ (7)
@@
wherein we denote the sample average of $x_{i}$ for men and women as $\bar{x}_{m}$ and $\bar{x}_{w}$, respectively, and we denote male and female coefficients on $x_{i}$ as $\beta_{m,x}$ and $\beta_{w,x}$, respectively. [^14] Fairlie’s (fairlie2005extension) decomposition for nonlinear models considers the case in which the marginal distribution of $x_{i}$ for women is set equal to men’s, while ours considers the case in which the mean of $x_{i}$ for women is set equal to men’s via a level shift of women’s distribution. We view the counterfactual we consider as likely simpler to understand and simpler to implement for policymakers, though describing $\beta_{w,x}$ as the “effect” of changing $x_{i}$ for women requires assumptions beyond those that either we or fairlie2005extension make. We similarly estimate the contribution to the average gender gap of $x_{i}$’s differential effect on $Y_{i}$ between men and women as
$\hat{\Delta}_{\beta_{x}}=\sum_{i=1}^{N_{w}}\frac{\Phi(X_{i}\hat{\beta}_{w}+x_{i}(\hat{\beta}_{m,x}-\hat{\beta}_{w,x}))-\Phi(X_{i}\hat{\beta}_{w})}{N_{w}}.$ (8)
@@
In addition to estimating the contribution of each variable and its coefficient on the gender-intention gap, we also estimate the contribution of the entire vector $X_{i}$ on the gap as
$\hat{\Delta}_{X}=\sum_{i=1}^{N_{w}}\frac{\Phi(X_{i}\hat{\beta}_{w}+(\bar{X}_{m}-\bar{x}_{w})\hat{\beta}_{w})-\Phi(X_{i}\hat{\beta}_{w})}{N_{w}},$ (9)
@@
defining the vector of the means of all elements in $X$ for people with $W_{i}=w$ as $\bar{X}_{w}$. Similarly, we estimate the contribution of the entire vector of gender differences in coefficients on the gender-intention gap as
$\hat{\Delta}_{\beta}=\sum_{i=1}^{N_{w}}\frac{\Phi(X_{i}\hat{\beta}_{w}+X_{i}(\hat{\beta}_{m}-\hat{\beta}_{w}))-\Phi(X_{i}\hat{\beta}_{w})}{N_{w}}.$ (10)
@@
$\hat{\Delta}_{R}=\sum_{i=N_{w}+1}^{N}\frac{\Phi(X_{i}\hat{\beta}_{m})}{N_{m}}-\sum_{i=1}^{N_{w}}\frac{\Phi(X_{i}\hat{\beta}_{w})}{N_{w}}-\hat{\Delta}_{X}-\hat{\Delta}_{\beta}.$ (11)
Table 2 contains estimated means of the variable or coefficient listed in each row for men, women, and the contribution of that item to the intention gap, using predicted values from a naive probit in columns 1–3, and using our proposed selection correction model in columns 4–6. $X$ is observed for all individuals regardless of survey response, so the means of $X$ by gender are the same for the uncorrected model and the corrected model. Columns 3 and 6 report estimated contributions of each row quantity to the gender gap as described in equations (7), (8), (9), (10), and (11). We perform our decomposition for all variables in Table 1, with term fixed effects included in $X_{i}$ contributing only to the unexplained gap.

### v4

@@ Abstract
When subjects who respond to requests for data, such as in surveys or post-treatment follow-up, are not representative of the population as a whole, inferences drawn from the data can be misleading. We show that if subjects’ accumulated requests and responses over time are recorded and organized as panel data, requests can be used as instruments to correct for nonresponse bias even if total requests are not randomized between subjects. We demonstrate our method by estimating an 18-percentage-point gender gap in entrepreneurial career intentions using a survey of undergraduates at the University of Wisconsin-Madison.
@@ 1 Introduction
In Section 3, we demonstrate the use of nonrandom request instruments by estimating the gender gap in entrepreneurial career intentions among undergraduates using survey data from the University of Wisconsin-Madison. Estimating a parametric van1981demand model for binary outcomes, we find an 18-percentage-point nonresponse-corrected male-female gap in entrepreneurial intentions. We find statistically insignificant evidence of positive selection for men and negative selection for women driving the 20 percentage point uncorrected gap among respondents. This model is identified with two requests, so we also implement and pass a “predictable trends” overidentification test of the null that later requests have no direct effect on the outcome, supporting the validity of our parametric and nonparametric assumptions in this application. In Supplemental Appendix B, we estimate labor market outcomes using a synthetic version of the Norway in Corona Times survey constructed to match estimates reported by dutz2021selection. The 95% confidence intervals of our estimated population means cover ground truth averages from administrative data reported by these authors for five out of six survey variables, with respondent averages differing significantly from the ground truth for all variables.
@@ 2.2 Modeling Nonresponse
We assume that observed variables are realizations of potential outcomes, following rubin1974estimating. First, let $S_{it}=S_{it}(R_{it})\prod_{j=0}^{t-1}(1-S_{ij}(R_{ij}))$ where $S_{it}(r)\in\{0,1\}$ indicates subject $i$’s willingness to respond to $r$ requests at time tt. [^7] We use this definition in lieu of $S_{it}=S_{it}(R_{it})$ because we assume that subjects respond at most once. Similarly, let $Y_{it}=Y_{it}(S_{it},R_{it})=Y_{it}(1,R_{it})S_{it}+Y_{it}(0,R_{it})(1-S_{it})$ with $Y_{it}(1,r)$ denoting the response subject $i$ would give at time $t$ if were to respond given $r$ accumulated requests, with $Y_{it}(0,r)$ missing for all $i$, $t$, and $r$. $S_{it}(r)$ and $Y_{it}(1,r)$ are defined for all $r\in\{0,...,\max(R_{it})\}$, respectively, while they are observed only for the realized $R_{it}$.
@@
It follows that we can learn about $m(u)$ if we can estimate $\mathbb{E}[Y_{i}^{*}|S_{it}(r)-S_{it}(r^{\prime})=1]$, which we refer to as the local average response for $r-r^{\prime}$ compliers. If $m(u)$ is parameterized, as in a heckman1979sample model, point estimation is possible if there are weakly more identified local average responses than parameters. Otherwise, $m(u)$ can be bounded using methods such as those of horowitz2000nonparametric or lee2009training. Local average responses are identified as
@@ 2.3 Local Average Response Identification Assumptions
This assumption strengthens the independence assumption of imbens1994identification by disallowing requests from being selectively sent to subjects with nonrepresentative potential responses to that request or any prior request.Randomizing requests and setting $T=1$, as proposed by dinardo2021practical, supports this assumption, with it also holding for $T>1$ with uniform requests if potential responses and response choices are time-invariant on average. Less extreme coarsening of the time index, for instance defining $t$ in weeks instead of minutes, can support this assumption by essentially averaging over temporal idiosyncrasies that occur in continuous time.
@@
The importance of avoiding experimenter demand effects or priming in surveys is well established (stantcheva2023run). Our assumption expresses this as a special case of similar assumptions made in the context of sample selection by heckman1979sample and in the context of treatment effect identification by angrist1996identification. In contrast with intertemporal independence, exclusion is more likely to be violated due to continuous time variation in potential responses and requests if time periods are defined coarsely, as subjects receiving different numbers of requests in the same discrete time period may receive them at different points in continuous time.
@@
This assumption is weaker than $Y_{it}(1,r)=Y_{i}^{*}$, which is common in the related literature, because it allows for idiosyncratic shocks that cancel out within complier groups, such as if students’ potential responses regarding career intentions change over time in response to new information, such as grades on assignments.Intertemporal independence and exclusion rule out systematic time-varying measurement error and measurement error caused by requests, respectively, but they do not rule out systematic inaccuracies within or across complier groups. For example, if less conscientious subjects give systematically inaccurate responses to late requests, nonresponse corrections using request instruments will tend to replicate similar inaccuracy in their predictions for non-respondents.
@@ 3 Application: Gender Entrepreneurship Gap
We use our method to estimate the gender gap in entrepreneurial aspirations among undergraduate students. There is a large gender gap in entrepreneurship among working-age adults (aldrich2005entrepreneurship), perhaps driven at least in part by a gender gap in early-stage funding for new ventures (canning2012women; greene2003women). Our paper contributes to this literature by investigating pre-market gender gaps in entrepreneurial career interest, which may be more likely to develop due to intrinsic interest gaps or pre-market discrimination rather than labor market discrimination. This application showcases the importance of researcher judgments in defining the time index, adjusting for nonrandom variation in requests, and testing modeling assumptions in cases where there are more requests than are needed for identification. In Supplemental Appendix B, we demonstrate our method’s accuracy by showing that nonresponse-corrected estimates are closer to ground truth values from administrative data than respondent averages for the Norway in Corona Times Survey, using survey and ground truth averages reported by dutz2021selection.
@@ 3.1 Data
We use a survey of the entrepreneurial intentions of the undergraduate population at University of Wisconsin–Madison that was implemented every fall from 2015 to 2022, as well as in spring 2020 and 2021. Students could choose between “Yes,” “No,” or “I Don’t Know,” when responding to a question regarding whether they intend to pursue a career in entrepreneurship. We code an “Intention” variable equal to 1 if the student answered “Yes” and 0 if they answered “No” or “I Don’t Know.” We observe term-specific survey request stratum indicators for groups of students who are sent requests at the same time through survey software. We observe timestamps for each request made to each stratum and for each student’s response if they respond. Students who opted out of the survey in a prior year have no stratum indicators and no request or response timestamps in the raw data, whereas students who respond to early requests have stratum indicators and request timestamps for all intended requests in raw data, though they did not receive additional actual requests after responding.
@@ 3.2 Constructing a Request-Response Panel
Our results from Section 2 involve panel data wherein subjects are observed multiple times as data requests are made, so we construct such a panel for each term. We index terms/semesters by $s$, with retained responses, retained response choices, and accumulated requests denoted by $\hat{Y}_{ist}$, $\hat{S}_{ist}$, and $R_{ist}$, respectively. In our setting, students who opted out of the survey in a prior term and students who respond early in a given term do not receive subsequent requests. It follows that actual requests almost surely violate the independence assumption in Section 2, as students receiving each number of requests differ systematically. We address this by randomly assigning opt-outs to strata ex post and assuming $S_{ist}(r)=0$ for all $t$ and all $r$ for such students, then imputing intended request timestamps for each student using the request timestamps of nonrespondents in their stratum.
For each student-term in a stratum, we produce observations for $t$ in $0,1,...,R_{i}$ such that time periods are defined as intervals in which a subject has received a given number of intended requests, with $R_{ist}=t$ for all $i$, $s$, and $t$. Independence and monotonicity require that the times at which requests were made had representative potential responses and that subjects had sufficient time to respond between requests, respectively. We report request timestamps in Appendix A, revealing that the shortest interval between requests is three days, which we assume was adequate time for students to respond. We set $S_{ist}=1$ if subject $i$’s response timestamp is between those defining the beginning and end of period $t$, with $Y_{ist}$ given by their response to the survey if $S_{ist}=1$. We forward-impute retained response choices $\hat{S}_{ist}$ and retained responses $\hat{Y}_{ist}$ as described in Section 2 for each value of $s$.
@@ 3.3 Nonresponse-Corrected Entrepreneurial Intentions by Gender
We begin by estimating the raw gender gap in entrepreneurial intentions without adjusting for the covariates listed in Table 1. Because requests vary in number from two to seven between terms, we drop all observations with $R_{it}>2$ to support our independence assumption for the current specification. We estimate a parametric van1981demand selection model via maximum likelihood, separately for men and women. Specifically, we assume
$\hat{S}_{ist}=\begin{cases}&1\mbox{ if }Z_{ist}\alpha+u_{ist}\geq 0,\\ &0\mbox{ otherwise,}\end{cases}$ (4)
$\hat{Y}_{ist}=\begin{cases}1\mbox{ if }X_{is}\beta+\epsilon_{ist}\geq 0\quad\&\quad&\hat{S}_{ist}=1,\\ 0\mbox{ if }X_{is}\beta+\epsilon_{ist}<0\quad\&\quad&\hat{S}_{ist}=1,\\ 0\mbox{ if }&\hat{S}_{ist}=0,\end{cases}$ (5)
@@
where $X_{is}=1$, and $Z_{ist}=[X_{is},I(R_{ist}=2)]$. We assume $(\epsilon_{ist},u_{ist})$ are independent and identically distributed between $i$, with potential dependence allowed between $s$ and $t$ for each $i$. $\rho$ gives the correlation between outcome residuals $\epsilon_{ist}$ and response preference $u_{ist}=\Phi^{-1}(1-U_{ist})$, where $U_{ist}$ is $i$’s response aversion percentile in the notation of Section 2 and $\Phi^{-1}(\cdot)$ denotes the inverse of the normal CDF. [^9] Our 15% response rate is insufficient for informative bounds from conservative approaches such as that of horowitz2000nonparametric, and we doubt that gender satisfies the monotonicity assumption of lee2009training or behaghel2015please, rendering the “local gender gap” among survey respondents unidentified, to say nothing of its policy-relevance.
This gender-specific constant-only model implies that expected intention as a function of response aversion is $m(u)=\Phi((\beta+\rho\Phi^{-1}(1-u))/\sqrt{1-\rho^{2}})$, with the unconditional expectation expressed as $\int_{0}^{1}m(u)du=\Phi(\beta)$. We graphically present our estimates of $m(u)$ for men and women in Figure 1 for $u\in[0,1]$. Local average responses and sample averages for compliers to the first two requests are shown for reference. The levels of entrepreneurial intention for early and late respondents show a decreasing pattern for men and an increasing pattern for women, as in Table 1. Differences in selection bias between men and women, reflected by their $\rho$ estimates, explain the difference between the 20 percentage point uncorrected gap estimate and the 12 percentage point corrected gap estimate.
@@ 3.4 Oaxaca-Blinder-Kitagawa Decomposition
To implement the decomposition, we estimate the model in (4)-(6) jointly for men and women via maximum likelihood using gender-specific parameters $(\beta_{m},\alpha_{m},\rho_{m})$ for men and $(\beta_{w},\alpha_{w},\rho_{w})$ for women. To avoid overweighting terms in which more requests were sent, we weight each observation by the inverse of the total number of requests sent to that student in that term. We define $X_{is}$ as all covariates listed in Table 1 along with term fixed effects, and we define $Z_{ist}$ as binary indicators for each value in $R_{ist}=1,2...,7$ interacted with $X_{is}$.101010blandhol2022tsls recommend rich covariate interactions for instrumental variables when estimating local average treatment effects via two stage least squares. The estimator and target parameters are different here, but we find their argument convincing that if model assumptions hold conditional on $X_{i}$, then $X_{i}$ should be flexibly interacted with instruments to support identification. We order the sample of $N_{w}$ women and $N_{m}$ men such that subjects with $i=1,...,N_{w}$ are women and those with $i=N_{w}+1,...,N$ are men, with $N=N_{w}+N_{m}$.
@@
Notes: Uncorrected estimates obtained by estimating a van1981demand model using only the final request observation for each student-term with the constraint $\rho=0$. Corrected estimates obtained by estimating the same model using all student-term-request observations with unconstrained $\rho$, with observations weighted by the inverse of total requests by student-term. Group-specific population mean estimates of variables and parameters, as indicated, are in columns (1), (2), (4), and (5), with sample means of $X$ calculated across student-term observations. Columns (3) and (6) give uncorrected and corrected estimated contributions of each variable or coefficient to the estimated gap in entrepreneurial intentions, using equations (7), (8), (9), (10), and (11). $\mathbb{E}[\mathbb{E}[Y_{is}^{*}|S_{isT}=1,X_{i}]]$ is estimated as the average of $\Phi(X_{is}\hat{\beta}^{uncorrected})$ and $\mathbb{E}[\mathbb{E}[Y_{is}^{*}|X_{is}]]$ is estimated as the average of $\Phi(X_{is}\hat{\beta}^{corrected})$ over student-term observations. Term fixed effects, included in $X_{i}$ to support the independence assumption, are treated as part of the unexplained gap. Standard errors clustered at the individual level are in parentheses.
We estimate the contribution of $x_{is}\subset X_{is}$ to the entrepreneurial intention gap as
$\hat{\Delta}_{x}=\sum_{i=1}^{N_{w}}\frac{\Phi(X_{is}\hat{\beta}_{w}+(\bar{x}_{m}-\bar{x}_{w})\hat{\beta}_{w,x})-\Phi(X_{is}\hat{\beta}_{w})}{N_{w}},$ (7)
@@
wherein we denote the sample average of $x_{i}$ for men and women as $\bar{x}_{m}$ and $\bar{x}_{w}$, respectively, and we denote male and female coefficients on $x_{i}$ as $\beta_{m,x}$ and $\beta_{w,x}$, respectively. [^11] Fairlie’s (fairlie2005extension) decomposition for nonlinear models considers the case in which the marginal distribution of $x_{is}$ for women is set equal to men’s, while ours considers the case in which the mean of $x_{is}$ for women is set equal to men’s via a level shift of women’s distribution. We view the counterfactual we consider as likely simpler to understand and simpler to implement for policymakers, though describing $\beta_{w,x}$ as the “effect” of changing $x_{is}$ for women requires assumptions beyond those that either we or fairlie2005extension make. We similarly estimate the contribution to the average gender gap of $x_{is}$’s differential effect on $Y^{*}_{is}$ between men and women as
$\hat{\Delta}_{\beta_{x}}=\sum_{i=1}^{N_{w}}\frac{\Phi(X_{is}\hat{\beta}_{w}+x_{is}(\hat{\beta}_{m,x}-\hat{\beta}_{w,x}))-\Phi(X_{is}\hat{\beta}_{w})}{N_{w}}.$ (8)
@@
In addition to estimating the contribution of each variable and its coefficient on the gender-intention gap, we also estimate the contribution of the entire vector $X_{is}$ on the gap as
$\hat{\Delta}_{X}=\sum_{i=1}^{N_{w}}\frac{\Phi(X_{is}\hat{\beta}_{w}+(\bar{X}_{m}-\bar{X}_{w})\hat{\beta}_{w})-\Phi(X_{is}\hat{\beta}_{w})}{N_{w}},$ (9)
@@
defining the vector of sample means of all covariates in $X_{is}$ for men as $\bar{X}_{m}$ and for all women as $\bar{X}_{w}$. Similarly, we estimate the contribution of the entire vector of gender differences in coefficients on the gender-intention gap as
$\hat{\Delta}_{\beta}=\sum_{i=1}^{N_{w}}\frac{\Phi(X_{is}\hat{\beta}_{w}+X_{is}(\hat{\beta}_{m}-\hat{\beta}_{w}))-\Phi(X_{is}\hat{\beta}_{w})}{N_{w}}.$ (10)
@@
$\hat{\Delta}_{R}=\sum_{i=N_{w}+1}^{N}\frac{\Phi(X_{is}\hat{\beta}_{m})}{N_{m}}-\sum_{i=1}^{N_{w}}\frac{\Phi(X_{is}\hat{\beta}_{w})}{N_{w}}-\hat{\Delta}_{X}-\hat{\Delta}_{\beta}.$ (11)
Table 2 contains estimated means of the variable or coefficient listed in each row for men, women, and the contribution of that item to the intention gap, using predicted values from a naive probit in columns 1–3, and using our proposed selection correction model in columns 4–6. $X_{is}$ is observed for all individuals regardless of survey response, so the means of $X_{is}$ by gender are the same for the uncorrected model and the corrected model. Columns 3 and 6 report estimated contributions of each row quantity to the gender gap as described in equations (7), (8), (9), (10), and (11). We perform our decomposition for all variables in Table 1, with term fixed effects included in $X_{is}$ contributing only to the unexplained gap.